# Energetic Electrons Accelerated and Trapped in a Magnetic Bottle above a Solar Flare Arcade

### v1

@@ 4 Discussion and Conclusion
The strong concentration of nonthermal electrons within the magnetic bottle region has important implications for electron acceleration and transport processes. First, it is highly likely that the nonthermal electrons are primarily accelerated in the magnetic bottle region itself rather than injected from elsewhere, such as high up in the reconnection current sheet or low in the loop legs. As shown in Figure 8(d), our joint observation–modeling results suggest that the nonthermal electron density rapidly drops to $<$1% of the peak value outside the ALT magnetic bottle region in all directions. Such a large density contrast implies that if the injection-and-trapping process is responsible, it must be a slow process. For instance, assuming all the electrons are injected from the upper current sheet, pure injection of 50 keV electrons without any loss takes $\tau^{\rm inj}\approx N_{e}^{\rm ALT}/F_{e}^{\rm inj}=n_{e}^{\rm ALT}L_{y}^{\rm ALT}A_{xz}^{\rm ALT}/(n_{e}^{\rm CS}v^{\rm inj}A_{xz}^{\rm CS})\approx(n_{e}^{\rm ALT}/n_{e}^{\rm CS})(A_{xz}^{\rm ALT}/A_{xz}^{\rm CS})L_{y}^{\rm ALT}/v^{\rm inj}$, where $n_{e}^{\rm ALT}/n_{e}^{\rm CS}>100$ is the ratio of the nonthermal electron density between the ALT and current sheet region, $A_{xz}^{\rm ALT}/A_{xz}^{\rm CS}>50$ is the expansion factor of the cross-section between the magnetic bottle region (with a width of $L_{x}^{\rm ALT}\approx 10$ Mm) and the narrow current sheet (with a width of $L_{x}^{\rm CS}\approx 0.2$ Mm, the grid size of the numerical model), $L_{y}^{\rm ALT}$ is the vertical extension of the ALT region taken to be $\sim$5 Mm as suggested in the model, and $v^{\rm inj}$ is the injection speed of the electrons, taken the kinetic speed of 50 keV electrons $v_{e}^{50}\approx 0.41c$ as the upper limit. The estimated injection timescale to produce the ALT source is $\tau_{\rm inj}>200$ s, even without any loss. This timescale is much longer than the acceleration timescale inferred from HXR and radio emissions, which can display rapid fluctuations at second or even subsecond scales (e.g., Fletcher et al. 2011 and references therein), . While we cannot entirely rule out such an injection-and-trapping scenario owing to the insufficient observational constraints for the precise dimensions of the ALT region (particularly in the line-of-sight $z$ direction), we conclude that local acceleration and trapping in this region is a more likely scenario to account for the profound concentration of the nonthermal electrons in the ALT magnetic bottle region.
Finally, our results provide new insights into understanding the puzzling departure from equipartition between the upward-escaped and downward-retained energetic electron population reported by studies that combine remote-sensing and in situ observations, which have concluded consistently that only 0.1–1% of the flare-accelerated energetic electrons manage to escape to interplanetary space (Lin, 1974; Krucker et al., 2007; Wang et al., 2021, 2023a). Such a large imbalance not only poses a challenge for understanding the particle acceleration and transport processes, but also has important implications for space weather in solar and systems. In our scenario, this imbalance naturally occurs because the primary acceleration site is the magnetic bottle located below the X point, with the newly reconnected, cusp-shaped field lines in the large-scale current sheet acting as a nearly closed structure to limit the upward-directed electrons from escaping (Figure 10). Along with efficient trapping, as in our case, the energetic electron density reaching the X point can be only $<$1% of that at the core acceleration region, giving rise to an extremely small fraction of these energetic electrons that enter interplanetary space.
Moreover, in order to achieve a more definitive understanding, extensive studies are needed for a large sample of flare events with different intensities and geometries. Last but not least, to make further progress, next-generation telescopes capable of performing radio and HXR imaging spectroscopy with orders of magnitude higher dynamic range and sensitivity, as well as more sophisticated models are desired. Notable future telescope concepts include the Frequency Agile Solar Radiotelescope (FASR; Gary et al. 2023 and Chen et al. 2023) in radio wavelengths and the Focusing Optics X-ray Solar Imager (FOXSI; Christe et al. 2016) or its variants (e.g., COMPLETE; Caspi et al. 2023) in X-rays.
This work is primarily funded by NASA Heliophysics Supporting Research grant 80NSSC20K1318 to NJIT, SAO, UMN, and LANL. B.C. and S.Y. receive additional support from NSF grants AST-1735405 and AST-2108853 to NJIT. The authors are grateful to Dr. Harry Warren for providing the Hinode/EIS temperature and emission measure maps. They also thank Drs. Alexey Kuznetsov and Gregory Fleishman for making their fast gyrosynchrotron codes available. X.K. is supported by NSFC grants 42074203 and 11873036. C.S. is supported by NASA grants 80NSSC19K0853, 80NSSC21K2044, and NSF AST 2108438 to SAO. X.L. acknowledges the support from NASA through grant 80NSSC21K1313, NSF grant AST-2107745, SAO through subcontract SV1-21012, and LANL through subcontract No. 622828. The Expanded Owens Valley Solar Array (EOVSA) was designed and built and is now operated by the New Jersey Institute of Technology (NJIT) as a community facility. EOVSA operations are supported by NSF grant AGS-2130832 and NASA grant 80NSSC20K0026 to NJIT. The authors are grateful to the RHESSI, SDO, SOHO/LASCO, and GOES-R/SUVI teams for making their data publicly available.
Appendix A X-ray data analysis
Figure 11: RHESSI X-ray images for the time of interest on 2017 September 10. The time interval used for imaging is 16:10:00-16:11:10 UT. The energy ranges used for imaging are 10–16 keV, 16–25 keV, and 25–60 keV, shown in panels (a), (b), and (c) respectively. The solid and dashed lines denote the 50% and 10% contours. Panel (d) shows X-ray images of all three energy ranges overlaid on SDO/AIA 131 Å (Fe XXI) EUV image at the closest time (shown in reversed gray-scale).
X-ray spectral analysis is performed using the ospex software (Schwartz et al., 2002). We use the data from detector 3 (which had the lowest sensitivity at the time and hence was less affected by the pile-up effect) to obtain the X-ray count flux spectrum between 16:10:00 UT and 16:11:08 UT. The resulting X-ray photon flux spectrum after applying the instrument response matrix is shown in Figure 4 as black curves. Spectral modeling is performed in the 12–60 keV range using two isothermal functions and one nonthermal bremsstrahlung model. Also included in the spectral modeling is the pile-up module. We performed spectral analysis assuming both the thick- and thin-target regimes. In both regimes, the thermal components, which dominate the spectrum below $\sim$30 keV, include a hot $\sim$15 MK looptop source with a volume emission measure of $\approx 2\times 10^{51}\,\mathrm{cm}^{-3}$. A secondary superhot component with a temperature of 27–28 MK and a volume emission measure of $\approx 4\times 10^{49}\,\mathrm{cm}^{-3}$ is also present. Assuming a source volume $V\approx(10\,\mathrm{Mm})^{3}$ according to the size of the source shown in the X-ray image, the plasma density for the two thermal components becomes $\approx 1\times 10^{12}$ cm-3 and $2\times 10^{11}$ cm-3, respectively. Such a high density associated with the hot 15 MK component is sufficient to stop all nonthermal electrons up to 70 keV in the source (with a half-width of $\sim$5 Mm) through Coulomb collisions (Tandberg-Hanssen & Emslie, 1988).
@@ Appendix A X-ray data analysis
As RHESSI was approaching the end of its operations at the time of the observation, only 4 of the 9 detectors operated nominally, which limited its imaging capabilities. Furthermore, SOL2017-09-10 was one of the brightest flares ever observed by RHESSI. Despite that both attenuators were inserted, the count rate nevertheless stayed high, and pileup effects occurred (i.e., two photons arrive essentially simultaneous, and they are therefore measured as a single photon with the summed energy of the two individual photons). While the time used for the analysis is $\sim$12 minutes past the flare peak, which has lessened the pileup issue, and pileup correction following the standard procedure has been applied in our fitting, the spectral analysis is possibly still pushing to the limits of the instrumentation. Hence, despite the fact that the spectral fitting results are consistent with other complementary data, we suggest that they should be considered with some level of caution. For imaging, pileup correction is not available, making the exact partition of the X-ray photons between the looptop and ALT X-ray source undetermined. However, as pileup only affects imaging by removing a rather small fraction of photons as they pile up and appear at higher energies, the imaging morphology is not much affected for energies below $\sim$36 keV (two times the peak of the count spectrum at $\sim$18 keV). We therefore conclude that the double-source structure seen in the X-ray images is trustworthy. However, we can not draw any firm conclusion on the true shape and brightness of the ALT X-ray source.
Appendix B Microwave data analysis
EOVSA observed the SOL2017-09-10 X8.2 flare event in the range 2.5–18 GHz with 31 evenly spaced spectral windows, each of which has a bandwidth of 160 MHz. The center frequencies of these spectral windows are $\nu=2.92+0.5n$ GHz, where $n$ is the spectral window number from 0 to 30. Different from Chen et al. (2020b) and Fleishman et al. (2022), which analyzed the earlier and main impulsive phase of the event, respectively, the time of interest for this study is around 16:10 UT during the gradual phase of the flare when the magnetic flux rope has already propagated to a large distance, allowing detailed observation–modeling comparison. Methods used for calibrating and imaging the EOVSA data are identical to our earlier studies (see Gary et al. 2018 and Chen et al. 2020b). The integration time used for synthesis imaging is 4 s, and a circular beam with an FWHM size of $10^{\prime\prime}.2/[\nu/{10\ \rm GHz}]$ is used to restore the final microwave images after deconvolution. Figure 5 shows an example of multi-frequency EOVSA images from 3.4 GHz ($n=1$) to 10.4 GHz ($n=15$) at 16:10:36 UT.
Each spatially resolved microwave spectrum used for spectral analysis is derived from the average brightness temperature in an $8^{\prime\prime}\times 8^{\prime\prime}$ area (corresponds to the resolution at $\sim$12.4 GHz). For each spatially resolved spectrum, a power-law nonthermal electron distribution $dn_{e}(\varepsilon)/d\varepsilon$ with a spectral index of $\delta^{\prime}$ and total nonthermal density $n_{\rm nth}^{>50}$ above 50 keV is used to model the spectrum based on the fast gyrosynchrotron codes (Fleishman & Kuznetsov, 2010). Other model parameters used in the fit include the magnetic field strength $B$, thermal plasma density $n_{\rm th}$ of the source, and the viewing angle $\theta$ with respect to the magnetic field direction. The low- and high-energy cutoff of the electron distribution $\varepsilon_{\rm min}$, $\varepsilon_{\rm max}$, and the plasma temperature are fixed to 25 keV, 10 MeV, and 15 MK, respectively. For the Markov chain Monte Carlo analysis, similar to Chen et al. (2020b), we use a logarithmic likelihood function in the following form:
$\ln p=-\frac{1}{2}\sum_{n}\left[(T_{b,i}^{\rm o}-T_{b,i}^{\rm m})^{2}/\sigma_{T_{b,i}}^{2}+\ln(2\pi\sigma_{T_{b,i}}^{2})\right],$ (B1)
where $T_{b,i}^{\rm o}$ and $T_{b,i}^{\rm m}$ are the observed and modeled brightness temperature at frequency $\nu_{i}$, respectively, and $\sigma_{T_{b,i}}$ is the corresponding uncertainty estimated by adding the r.m.s. brightness temperature of a region in the image without any sources and a fractional error in the source brightness temperature (assumed to be 15%) in quadrature. We then use emcee (Foreman-Mackey et al., 2013), a Python implementation of the affine-invariant MCMC ensemble sampler (Goodman & Weare, 2010), to sample the parameter space according to the likelihood function. The multi-parameter posterior distribution allows us to find the best-fit model parameter values, taken as the median value of the samples in each marginalized distribution (i.e., the 50% percentile). The lower and upper 1-$\sigma$ uncertainties are taken as the 16% and 84% percentile of the samples in each distribution, respectively. For the MCMC analysis of each spectrum, we use 100 “walkers” to sample the parameter space, each of which draws a total of 8000 samples.
Appendix C MHD and Particle Modeling
@@
Figure 12: Data-informed MHD and particle modeling of the 2017 September 10 flare during its gradual phase at 16:10 UT. (a) and (b) Plasma density and temperature maps derived from Hinode/EIS measurements of the Fe XXIV line pairs. (c) and (d) Hinode/EIS Fe Fe XXIV and SDO/AIA 131 Å (Fe XXI) intensity maps, which have a peak sensitivity to 18 MK and 10 MK plasma, respectively. The red contours outline the regions where the EIS intensity is saturated and has unreliable temperature/density diagnostics. Also, note the “X”-shaped artificial diffraction pattern in EIS and AIA images. (e)–(h) Distribution of plasma density $n_{\rm th}$, temperature $T$, magnetic field strength $B$, and vertical component of the plasma velocity $v_{y}$ from the MHD model. (i)–(l) Distribution of nonthermal electrons at 50 keV, 100 keV, 200 keV, and 500 keV, respectively, from the particle model. Black curves in all panels are magnetic field lines derived from the MHD model.
[…]
@@
The particle modeling adopts the method described in Kong et al. (2019). Similar approaches have been performed in several following studies (Kong et al., 2020, 2022b, 2022a; Li et al., 2022). We refer interested readers to these works for more detailed descriptions. Briefly, mono-energetic electrons of an initial energy of 20 keV are injected into the MHD model as pseudo particles with an isotropic angular distribution. The kinetic evolution of these particles in the simulation domain is modeled by solving the Parker transport equation, which takes the fluid velocity and magnetic field input from the MHD model. Particle transport in the magnetic field is mainly modulated by stochastic diffusion by well-developed turbulence with a Kolmogorov-type power spectrum. The construction of the diffusion coefficient follows the treatment in Giacalone & Jokipii (1999), with the perpendicular diffusion assumed to be 10% of the parallel diffusion. The simulation domain has an area of $102.0\ \text{Mm}\times 127.5\ \text{Mm}$ with a uniform grid size of 0.22 Mm. The output of the model, binned to a grid size of 1.275 Mm, is a spatial distribution of nonthermal electrons at different energies. In the model, the electron momentum is distributed evenly in logarithmic space, with a total of 40 samples between 20 keV and 5450 keV.
@@
Figures 12(i)–(l) shows the distribution of nonthermal electrons at four selected energies (50 keV, 100 keV, 200 keV, and 500 keV) from the particle model. Similar to the results in Kong et al. (2019), the nonthermal electrons at all energies are strongly concentrated in the magnetic bottle region. As discussed in the main text, we can draw an electron distribution function at each pixel of the model. Figure 8(c) shows examples from selected locations. All the electron spectra display a power-law shape in the $\sim$30–600 keV range, with those in the magnetic bottle region featuring a downward spectral break. Such broken power-law electron distributions in the model are formed by the combination of stochastic acceleration, trapping, and escaping processes (Kong et al., 2019; Li et al., 2022). The presence of a downward-breaking electron spectrum in the ALT magnetic bottle region is further supported by recent observations of the same flare during its early impulsive phase using combined EUV, X-ray, and microwave data (Chen et al., 2021).
Appendix D Emission Modeling
@@
Figure 13: Modeled energetic electron distribution and calculated hard X-ray and microwave spectra. (a) Distribution of $>$50 keV energetic electrons from the combined MHD and particle model. (b) Electron distribution from a sample pixel in the model (black box in (a)). Blue circles are the distribution from the model, and the black curve shows the broken power-law fit. (c) Calculated X-ray spectra based on the bremsstrahlung mechanism. Blue and black curves are those using direct integration of the discrete electron distribution and the broken power-law fit, respectively. (d) Similar to (c), but for the calculated microwave spectra based on gyrosynchrotron radiation.
@@
With maps of the thermal plasma properties and nonthermal electron distributions, one can calculate synthetic emission maps at various wavelengths and compare them with actual observations. To produce the synthetic SDO/AIA 131 Å EUV map, we only need the plasma density $n_{\rm th}$ and temperature $T$ from the 2.5-D MHD model. We first calculate the EUV intensity at each pixel with $I(x,y)=n_{\rm th}^{2}(x,y)G[T(x,y)]L_{z}$, where $G(T)$ is the temperature response function of the 131 Å band (O’Dwyer et al., 2010) and $L_{z}$ is the line of sight depth (fixed to be 10 Mm where flare-heated, $>8$ MK plasma is present and 1 Mm elsewhere to suppress the coronal background with an artificially high density as it is not optimized in the model), and then convolve the resulting image using a Gaussian kernel with a full width at half maximum (FWHM) width of $21^{\prime\prime}.$2 according to AIA’s point spread function (PSF). Although we do not have sufficient information on the third dimension (along the line of sight) and hence have to use a homogeneous assumption, the resulting EUV 131 Å map, shown in Figure 8(e), resembles the observations reasonably well (Figure 8(f)).
@@
In order to produce synthetic HXR and microwave maps, both the thermal plasma properties and nonthermal electron distributions are needed. For calculating synthetic HXR maps, at each pixel, we take the plasma density in the MHD model $n_{\rm th}(x,y)$ and the nonthermal electron distribution $dn_{e}(x,y,\varepsilon)/d\varepsilon$ and compute the X-ray photon flux $I_{X}(x,y,\epsilon)$ as a function of X-ray photon energy $\epsilon$ based on the thin-target bremsstrahlung theory. To perform the numerical calculations, we have adopted two commensurable approaches. One approach is to first fit the discrete nonthermal electron distribution from the particle model with a broken power-law form, and then supply the best-fit broken power-law function as the input to calculate the expected thin-target bremsstrahlung spectrum using existing tools available from the xray package within the sswIDL distribution [^2] Description of the software codes can be found at: https://hesperia.gsfc.nasa.gov/ssw/packages/xray/doc/brm_thin_doc.pdf. Another approach is, for every given X-ray photon energy $\epsilon$, we take the discrete model electron distribution and integrate the X-ray flux contributed by all energy bins numerically. Figure 13(b) shows an example electron distribution derived from a selected pixel in the magnetic bottle region. Blue symbols are the discrete distribution from the particle model, and the black solid line is the best-fit broken power-law function. The calculated X-ray photon spectra using direct integration and broken power-law fit of the electron distribution are shown in panel (c) as blue and black curves, respectively. The results show that they are in agreement with each other.
@@
For calculating the synthetic microwave spectrum from each pixel, we use the numerical codes developed by Kuznetsov & Fleishman (2021), which allows an input electron distribution in both the discrete numerical form and an analytical broken power-law form. Likewise, the results from the two different approaches are generally consistent with each other, although we found that the broken power-law approach gives better-defined microwave spectra as it effectively “smooths out” the occasional noise in the input electron distribution, especially in regions with low counting statistics. Therefore, we have adapted the broken power-law fit method to calculate the HXR and microwave spectra pixel-by-pixel, forming the spectrally resolved HXR and microwave maps. Finally, to compare with the observations, each image is convolved with a Gaussian function with the same FWHM width as the point spread function used to reconstruct the observed RHESSI and EOVSA images ($86^{\prime\prime}.$8 for RHESSI and $10^{\prime\prime}.2/[\nu/{10\ \rm GHz}]$ for EOVSA).

### v2

@@ 4 Discussion and Conclusion
As RHESSI was approaching the end of its operations at the time of the observation, only four of the nine detectors operated nominally, which limited its imaging capabilities. Furthermore, SOL2017-09-10 was one of the brightest flares ever observed by RHESSI. Despite both attenuators being inserted, the count rate nevertheless stayed high, and pileup effects occurred (i.e., two photons arrive essentially simultaneously, and they are therefore measured as a single photon with the summed energy of the two individual photons). While the time used for the analysis is $\sim$12 minutes past the flare peak, which has lessened the pileup issue, and pileup correction following the standard procedure has been applied in our fitting, the spectral analysis is possibly still pushing to the limits of the instrumentation. Hence, despite the fact that the spectral fitting results are consistent with other complementary data, we suggest that they should be considered with some level of caution. For imaging, pileup correction is not available, making the exact partition of the X-ray photons between the loop top and ALT X-ray source undetermined. However, as pileup only affects imaging by removing a rather small fraction of photons as they pile up and appear at higher energies, the imaging morphology is not much affected for energies below $\sim$36 keV (2 times the peak of the count spectrum at $\sim$18 keV). We therefore conclude that the double-source structure seen in the X-ray images is trustworthy. However, we cannot draw any firm conclusion on the true shape and brightness of the ALT X-ray source.
Appendix B: Microwave data analysis
EOVSA observed the SOL2017-09-10 X8.2 flare event in the range 2.5–18 GHz with 31 evenly spaced spectral windows, each of which has a bandwidth of 160 MHz. The center frequencies of these spectral windows are $\nu=2.92+0.5n$ GHz, where $n$ is the spectral window number from 0 to 30. Different from Chen et al. (2020b) and Fleishman et al. (2022), which analyzed the earlier and main impulsive phase of the event, respectively, the time of interest for this study is around 16:10 UT, during the gradual phase of the flare when the magnetic flux rope has already propagated to a large distance, allowing detailed observation–modeling comparison. Methods used for calibrating and imaging the EOVSA data are identical to our earlier studies (see Gary et al. 2018 and Chen et al. 2020b). The integration time used for synthesis imaging is 4 s, and a circular beam with an FWHM size of $10^{\prime\prime}.2/[\nu/{10\ \rm GHz}]$ is used to restore the final microwave images after deconvolution. Figure 5 shows an example of multifrequency EOVSA images from 3.4 GHz ($n=1$) to 10.4 GHz ($n=15$) at 16:10:36 UT.
Each spatially resolved microwave spectrum used for spectral analysis is derived from the average brightness temperature in an $8^{\prime\prime}\times 8^{\prime\prime}$ area (corresponding to the resolution at $\sim$12.4 GHz). For each spatially resolved spectrum, a power-law nonthermal electron distribution $dn_{e}(\varepsilon)/d\varepsilon$ with a spectral index of $\delta^{\prime}$ and total nonthermal density $n_{\rm nth}^{>50}$ above 50 keV is used to model the spectrum based on the fast gyrosynchrotron codes (Fleishman & Kuznetsov, 2010). Other model parameters used in the fit include the magnetic field strength $B$, thermal plasma density $n_{\rm th}$ of the source, and the viewing angle $\theta$ with respect to the magnetic field direction. The low- and high-energy cutoff of the electron distribution $\varepsilon_{\rm min}$, $\varepsilon_{\rm max}$, and the plasma temperature are fixed to 25 keV, 10 MeV, and 15 MK, respectively. For the MCMC analysis, similar to Chen et al. (2020b), we use a logarithmic likelihood function in the following form:
$\ln p=-\frac{1}{2}\sum_{n}\left[(T_{b,i}^{\rm o}-T_{b,i}^{\rm m})^{2}/\sigma_{T_{b,i}}^{2}+\ln(2\pi\sigma_{T_{b,i}}^{2})\right],$ (1)
where $T_{b,i}^{\rm o}$ and $T_{b,i}^{\rm m}$ are the observed and modeled brightness temperature at frequency $\nu_{i}$, respectively, and $\sigma_{T_{b,i}}$ is the corresponding uncertainty estimated by adding the rms brightness temperature of a region in the image without any sources and a fractional error in the source brightness temperature (assumed to be 15%) in quadrature. We then use emcee (Foreman-Mackey et al., 2013), a Python implementation of the affine-invariant MCMC ensemble sampler (Goodman & Weare, 2010), to sample the parameter space according to the likelihood function. The multiparameter posterior distribution allows us to find the best-fit model parameter values, taken as the median value of the samples in each marginalized distribution (i.e., the 50th percentile). The lower and upper 1-$\sigma$ uncertainties are taken as the 16th and 84th percentiles of the samples in each distribution, respectively. For the MCMC analysis of each spectrum, we use 100 “walkers” to sample the parameter space, each of which draws a total of 8000 samples.
Appendix C: MHD and Particle Modeling
@@
Figure 12: Data-informed MHD and particle modeling of the 2017 September 10 flare during its gradual phase at 16:10 UT. (a) and (b) Plasma density and temperature maps derived from Hinode/EIS measurements of the Fe XXIV line pairs. (c) and (d) Hinode/EIS Fe Fe XXIV and SDO/AIA 131 Å (Fe XXI) intensity maps, which have a peak sensitivity to 18 MK and 10 MK plasma, respectively. The red contours outline the regions where the EIS intensity is saturated and has unreliable temperature/density diagnostics. Also, note the “X”-shaped artificial diffraction pattern in the EIS and AIA images. (e)–(h) Distribution of plasma density $n_{\rm th}$, temperature $T$, magnetic field strength $B$, and vertical component of the plasma velocity $v_{y}$ from the MHD model. (i)–(l) Distribution of nonthermal electrons at 50 keV, 100 keV, 200 keV, and 500 keV, respectively, from the particle model. Black curves in all panels are magnetic field lines derived from the MHD model.
@@
The particle modeling adopts the method described in Kong et al. (2019). Similar approaches have been performed in several following studies (Kong et al., 2020, 2022b, 2022a; Li et al., 2022). We refer interested readers to these works for more detailed descriptions. Briefly, monoenergetic electrons of an initial energy of 20 keV are injected into the MHD model as pseudoparticles with an isotropic angular distribution. The kinetic evolution of these particles in the simulation domain is modeled by solving the Parker transport equation, which takes the fluid velocity and magnetic field input from the MHD model. Particle transport in the magnetic field is mainly modulated by stochastic diffusion by well-developed turbulence with a Kolmogorov-type power spectrum. The construction of the diffusion coefficient follows the treatment in Giacalone & Jokipii (1999), with the perpendicular diffusion assumed to be 10% of the parallel diffusion. The simulation domain has an area of $102.0\ \text{Mm}\times 127.5\ \text{Mm}$ with a uniform grid size of 0.22 Mm. The output of the model, binned to a grid size of 1.275 Mm, is a spatial distribution of nonthermal electrons at different energies. In the model, the electron momentum is distributed evenly in logarithmic space, with a total of 40 samples between 20 and 5450 keV.
@@
Figures 12(i)–(l) show the distribution of nonthermal electrons at four selected energies (50, 100, 200, and 500 keV) from the particle model. Similar to the results in Kong et al. (2019), the nonthermal electrons at all energies are strongly concentrated in the magnetic bottle region. As discussed in the main text, we can draw an electron distribution function at each pixel of the model. Figure 8(c) shows examples from four selected locations. All the electron spectra display a power-law shape in the $\sim$30–600 keV range, with those in the magnetic bottle region featuring a downward spectral break. Such broken power-law electron distributions in the model are formed by the combination of stochastic acceleration, trapping, and escaping processes (Kong et al., 2019; Li et al., 2022). The presence of a downward-breaking electron spectrum in the ALT magnetic bottle region is further supported by recent observations of the same flare during its early impulsive phase using combined EUV, X-ray, and microwave data (Chen et al., 2021).
Appendix D: Emission Modeling
@@
Figure 13: Modeled energetic electron distribution and calculated HXR and microwave spectra. (a) Distribution of $>$50 keV energetic electrons from the combined MHD and particle model. (b) Electron distribution from a sample pixel in the model (black box in (a)). Blue circles are the distribution from the model, and the black curve shows the broken power-law fit. (c) Calculated X-ray spectra based on the bremsstrahlung mechanism. Blue and black curves are those using direct integration of the discrete electron distribution and the broken power-law fit, respectively. (d) Similar to (c) but for the calculated microwave spectra based on gyrosynchrotron radiation.
@@
With maps of the thermal plasma properties and nonthermal electron distributions, one can calculate synthetic emission maps at various wavelengths and compare them with actual observations. To produce the synthetic SDO/AIA 131 Å EUV map, we only need the plasma density $n_{\rm th}$ and temperature $T$ from the 2.5D MHD model. We first calculate the EUV intensity at each pixel with $I(x,y)=n_{\rm th}^{2}(x,y)G[T(x,y)]L_{z}$, where $G(T)$ is the temperature response function of the 131 Å band (O’Dwyer et al., 2010) and $L_{z}$ is the LOS depth (fixed to be 10 Mm where flare-heated $>8$ MK plasma is present and 1 Mm elsewhere to suppress the coronal background with an artificially high density, as it is not optimized in the model), and then convolve the resulting image using a Gaussian kernel with an FWHM of $21^{\prime\prime}.$2 according to AIA’s point spread function (PSF). Although we do not have sufficient information on the third dimension (along the LOS) and hence have to use a homogeneous assumption, the resulting EUV 131 Å map, shown in Figure 8(e), resembles the observations reasonably well (Figure 8(f)).
@@
In order to produce synthetic HXR and microwave maps, both the thermal plasma properties and nonthermal electron distributions are needed. For calculating synthetic HXR maps, at each pixel, we take the plasma density in the MHD model $n_{\rm th}(x,y)$ and the nonthermal electron distribution $dn_{e}(x,y,\varepsilon)/d\varepsilon$ and compute the X-ray photon flux $I_{X}(x,y,\epsilon)$ as a function of X-ray photon energy $\epsilon$ based on the thin-target bremsstrahlung theory. To perform the numerical calculations, we have adopted two commensurable approaches. One approach is to first fit the discrete nonthermal electron distribution from the particle model with a broken power-law form, and then supply the best-fit broken power-law form as the input to calculate the expected thin-target bremsstrahlung spectrum using existing tools available from the xray package within the sswIDL distribution [^2] Description of the software codes can be found at https://hesperia.gsfc.nasa.gov/ssw/packages/xray/doc/brm_thin_doc.pdf. Another approach is, for every given X-ray photon energy $\epsilon$, we take the discrete model electron distribution and integrate the X-ray flux contributed by all energy bins numerically. Figure 13(b) shows an example electron distribution derived from a selected pixel in the magnetic bottle region. Blue symbols are the discrete distribution from the particle model, and the black solid line is the best-fit broken power-law function. The calculated X-ray photon spectra using direct integration and a broken power-law fit of the electron distribution are shown in panel (c) as blue and black curves, respectively. The results show that they are in agreement with each other.
@@
For calculating the synthetic microwave spectrum from each pixel, we use the numerical codes developed by Kuznetsov & Fleishman (2021), which allow an input electron distribution in both the discrete numerical form and an analytical broken power-law form. Likewise, the results from the two different approaches are generally consistent with each other, although we found that the broken power-law approach gives better-defined microwave spectra as it effectively “smooths out” the occasional noise in the input electron distribution, especially in regions with low counting statistics. Therefore, we have adapted the broken power-law fit method to calculate the HXR and microwave spectra pixel by pixel, forming the spectrally resolved HXR and microwave maps. Finally, to compare with the observations, each image is convolved with a Gaussian function with the same FWHM width as the point spread function used to reconstruct the observed RHESSI and EOVSA images ($86^{\prime\prime}.$8 for RHESSI and $10^{\prime\prime}.2/[\nu/{10\ \rm GHz}]$ for EOVSA).